### v1

@@ 2.1 Overview of FetalCLIP
FetalCLIP is a vision-language foundation model for fetal ultrasound analysis, developed by leveraging a large-scale dataset of paired fetal ultrasound images and captions (Fig. 1a). The dataset comprises 207,943 images from routine clinical scans encompassing 64 diverse clinician-labeled keywords across 6,493 patients with a mean gestational age of $148\pm 16$ days, supplemented by 2,092 image-caption pairs derived from a fetal ultrasound textbook [28] primarily focused on the fetal heart. As expert-provided image-level text descriptions were absent for the clinical dataset, we used GPT-4o to generate clinically sound captions based on gestational age, clinical labels, and pixel spacing, with details provided in the Methods. This approach ensured each image was paired with a unique, contextually relevant textual description. We adopted the contrastive language-image pretraining (CLIP) [11] framework to pretrain FetalCLIP by aligning images with their corresponding textual descriptions in a shared embedding space while minimizing similarity to unrelated pairs (Fig. 1b). The FetalCLIP architecture includes a ViT-L [29] image encoder, a Byte-Pair Encoding tokenizer [30], and a text encoder [31] capable of processing up to 117 tokens to accommodate the maximum token length in the rich clinical text descriptions. Our extensive evaluations demonstrated the superior performance of FetalCLIP over existing vision-language foundation models across various tasks in fetal ultrasound analysis (Fig. 1c), attributed to its pretraining on a large-scale dataset comprising predominantly clinical data with diverse fetal ultrasound keywords (Fig. 1d).
@@ 2.2 Zero-shot classification of standard fetal views
We conducted a study to evaluate FetalCLIP’s capability in classifying standard fetal ultrasound views using unseen data from different hospitals without any adaptation (Fig. 2a). FetalCLIP was compared against notable models in the field: (1) SonoNet [27], a specialized model specifically trained for classifying standard views of fetal ultrasound; (2) CLIP [11], a visual-language model (VLM) for natural images; and two visual-language foundation models tailored for the general medical domain, (3) BiomedCLIP [12] and (4) UniMed-CLIP [13]. We conducted the evaluation using the Planes DB [32] dataset, which consists of fetal ultrasound images acquired from two hospitals. The models were employed to classify five anatomical planes—abdomen, brain, cervix, femur, and thorax—as well as three subplanes within the brain—cerebellum, thalamic, and ventricular. Our experiment (Fig. 2b) demonstrated that FetalCLIP achieved an average F1 score of 87.1%, outperforming SonoNet by a notable margin of 17.2%, UniMed-CLIP by 37.6%, BiomedCLIP by 40.5%, and CLIP by 60.1%. The findings also highlight that CLIP lacks the contextual information required to differentiate between different fetal ultrasound anatomical views, and both BiomedCLIP and UniMed-CLIP struggle in distinguishing between brain subplanes. Our confusion matrix analysis (Extended Data Fig. 1) further added that SonoNet struggled with identifying the cervix view, while UniMed-CLIP could not distinguish the spine view from other fetal planes. In contrast, by harnessing large-scale fetal ultrasound data and incorporating semantic knowledge from textual descriptions during pretraining, FetalCLIP achieved remarkable accuracy in zero-shot classification of standard anatomical views.
@@ 4.1 FetalCLIP pretraining data curation
FetalCLIP was pretrained using two data sources, curated to develop a robust and generalizable visual-language foundation model for fetal ultrasound image analysis. The first data source consists of 207,943 fetal ultrasound images from Corniche Hospital Abu Dhabi, a facility specializing in Women and Newborn care. This data was supplemented by 2,092 image-caption pairs from a fetal ultrasound textbook [28] emphasizing fetal heart conditions, providing critically important clinical information for fetal health assessment in clinical practice.
@@ 4.1.1 Private hospital dataset
This data source constitutes the largest portion of the FetalCLIP pretraining dataset. The images were obtained from routine pregnancy scans at the hospital to monitor fetal development, often performed during the second trimester. During this stage, the fetus begins to resemble a child, and sonographers examine various anatomical planes to detect any potential health abnormalities. The average gestational age in this data source is $148\pm 16$ days, with 50% of the images lying between 20 weeks 0 days and 21 weeks 6 days. Some images were accompanied by clinicians’ text annotations embedded within the images, from which we identified 64 distinct keywords predominantly related to fetal anatomical structures (Fig. 1d). However, no associated medical report is available for each image. For this study, we utilized B-mode images and leveraged text annotations written by clinicians to serve as the basis for image labeling and text description generation for each image.
@@
For the heart images, clinicians provided subview annotations for LVOT, RVOT, 4-CH, and 3VV/3VT, providing fine-grained information. In contrast, 468 brain images (out of 5,297) were labeled by an expert into three brain subviews, i.e. thalamic, cerebellum, and ventricular. These expert-labeled brain images were then used to train a 5-fold CV brain subview classifier. Pseudo-labeling was applied to the remaining brain images, assigning subviews to images with probabilities exceeding 90%. Images with lower probabilities were categorized under the general ”Brain” class.
@@ 4.3 Zero-shot view classification
This zero-shot performance was evaluated using the Planes DB [32] dataset. The dataset contains six classes: abdomen, femur, brain, thorax, cervix, and a class labeled ”other,” representing diverse additional views. The brain category is further divided into three fine-grained subviews: cerebellum, thalamic, and ventricular. Each image was preprocessed by symmetrically padding with zeros to form a square and subsequently resizing to a uniform dimension of 224$\times$224. We first evaluated the performance on distinguishing five standard views, excluding the ”other” class, resulting in 8187 test images. To ensure a fair comparison with the specialist model (SonoNet), nine standard views were selected as target classes, i.e. abdomen, brain, femur, heart, kidney, lips & nose, profile, spine, and cervix. Secondly, we evaluated the model’s ability to classify the three brain subviews, resulting in 2949 test images. As SonoNet lacks dedicated classes for both cervix and thalamic, we mapped the ”other” class in SonoNet to cervix and thalamic for the first and second evaluations, respectively. In addition, the performance of the SonoNet models varied across sizes, with SN16, SN32, and SN64 achieving F1 scores of 67.5%, 69.9%, and 67.2%, respectively, where the mid-sized model (SN32) demonstrated the best performance.
@@ 5 Data Availability
This study involved two publicly available datasets for evaluating models on a subset of tasks, as well as private datasets for pretraining and additional evaluations. The public datasets were obtained from PlanesDB [32] and HC18 [33]. Due to regulations and privacy constraints, the private datasets cannot be released to the public.
@@ 7 Acknowledgements
We express our gratitude to Corniche Hospital in Abu Dhabi for providing prenatal scan data along with fetal heart scans, and to the Department of Health (DOH) Abu Dhabi for their support in approving the study which facilitates access to the anonymous data for internal purposes. We thank Alfred Z. Abuhamad for allowing us to leverage his book for foundation model pretraining.

### v2

@@ 2.1 Overview of FetalCLIP
FetalCLIP is a vision-language foundation model for fetal ultrasound analysis, developed by leveraging a large-scale dataset of paired fetal ultrasound images and captions (Fig. 1a). The dataset comprises 207,943 images from routine clinical scans encompassing 64 diverse clinician-labeled keywords across 6,493 patients with a mean gestational age of $148\pm 16$ days (21 weeks 1 day ± 2 weeks 2 days), supplemented by 2,092 image-caption pairs derived from a fetal ultrasound textbook [28] primarily focused on the fetal heart. As expert-provided image-level text descriptions were absent for the clinical dataset, we used GPT-4o to generate clinically sound captions based on gestational age, clinical labels, and pixel spacing, with details provided in the Methods. This approach ensured each image was paired with a unique, contextually relevant textual description. We adopted the contrastive language-image pretraining (CLIP) [11] framework to pretrain FetalCLIP by aligning images with their corresponding textual descriptions in a shared embedding space while minimizing similarity to unrelated pairs (Fig. 1b). The FetalCLIP architecture includes a ViT-L [29] image encoder, a Byte-Pair Encoding tokenizer [30], and a text encoder [31] capable of processing up to 117 tokens to accommodate the maximum token length in the rich clinical text descriptions. Our extensive evaluations demonstrated the superior performance of FetalCLIP over existing vision-language foundation models across various tasks in fetal ultrasound analysis (Fig. 1c), attributed to its pretraining on a large-scale dataset comprising predominantly clinical data with diverse fetal ultrasound keywords (Fig. 1d).
@@ 2.2 Zero-shot classification of standard fetal views
We conducted a study to evaluate FetalCLIP’s capability in classifying standard fetal ultrasound views using unseen data from different hospitals without any adaptation (Fig. 2a). FetalCLIP was compared against notable models in the field: (1) SonoNet [27], a specialized model specifically trained for classifying standard views of fetal ultrasound; (2) CLIP [11], a visual-language model (VLM) for natural images; and two visual-language foundation models tailored for the general medical domain, (3) BiomedCLIP [12] and (4) UniMed-CLIP [13]. We conducted the evaluation using the Planes DB [32] dataset, which consists of fetal ultrasound images acquired from two hospitals. The models were employed to classify five anatomical planes—abdomen, brain, cervix, femur, and thorax—as well as three subplanes within the brain—transcerebellum, transthalamic, and transventricular. Our experiment (Fig. 2b) demonstrated that FetalCLIP achieved an average F1 score of 87.1%, outperforming SonoNet by a notable margin of 17.2%, UniMed-CLIP by 37.6%, BiomedCLIP by 40.5%, and CLIP by 60.1%. The findings also highlight that CLIP lacks the contextual information required to differentiate between different fetal ultrasound anatomical views, and both BiomedCLIP and UniMed-CLIP struggle in distinguishing between brain subplanes. Our confusion matrix analysis (Extended Data Fig. 1) further added that SonoNet struggled with identifying the cervix view, while UniMed-CLIP could not distinguish the spine view from other fetal planes. In contrast, by harnessing large-scale fetal ultrasound data and incorporating semantic knowledge from textual descriptions during pretraining, FetalCLIP achieved remarkable accuracy in zero-shot classification of standard anatomical views.
@@ 4.1 FetalCLIP pretraining data curation
FetalCLIP was pretrained using two data sources, curated to develop a robust and generalizable visual-language foundation model for fetal ultrasound image analysis. The first data source consists of 207,943 fetal ultrasound images from Corniche Hospital Abu Dhabi, a tertiary-level referral unit providing expert care to women and neonates. This data was supplemented by 2,092 image-caption pairs from a fetal ultrasound textbook [28] emphasizing fetal heart conditions, providing critically important clinical information for fetal health assessment in clinical practice.
@@ 4.1.1 Private hospital dataset
This data source constitutes the largest portion of the FetalCLIP pretraining dataset. The images were obtained from routine pregnancy scans at the hospital to monitor fetal development, often performed during the second trimester. During this stage, the fetus begins to resemble a baby, and sonographers examine various anatomical planes to detect any potential health abnormalities. The average gestational age in this data source is $148\pm 16$ days, with 50% of the images lying between 20 weeks 0 days and 21 weeks 6 days. Some images were accompanied by clinicians’ text annotations embedded within the images, from which we identified 64 distinct keywords predominantly related to fetal anatomical structures (Fig. 1d). However, no associated medical report is available for each image. For this study, we utilized B-mode images and leveraged text annotations written by clinicians to serve as the basis for image labeling and text description generation for each image.
@@
For the heart images, clinicians provided subview annotations for LVOT, RVOT, 4-CH, and 3VV/3VT, providing fine-grained information. In contrast, 468 brain images (out of 5,297) were labeled by an expert into three brain subviews, i.e. transthalamic, transcerebellum, and transventricular. These expert-labeled brain images were then used to train a 5-fold CV brain subview classifier. Pseudo-labeling was applied to the remaining brain images, assigning subviews to images with probabilities exceeding 90%. Images with lower probabilities were categorized under the general ”Brain” class.
@@ 4.3 Zero-shot view classification
This zero-shot performance was evaluated using the Planes DB [32] dataset. The dataset contains six classes: abdomen, femur, brain, thorax, cervix, and a class labeled ”other,” representing diverse additional views. The brain category is further divided into three fine-grained subviews: transcerebellum, transthalamic, and transventricular. Each image was preprocessed by symmetrically padding with zeros to form a square and subsequently resizing to a uniform dimension of 224$\times$224. We first evaluated the performance on distinguishing five standard views, excluding the ”other” class, resulting in 8187 test images. To ensure a fair comparison with the specialist model (SonoNet), nine standard views were selected as target classes, i.e. abdomen, brain, femur, heart, kidney, lips & nose, profile, spine, and cervix. Secondly, we evaluated the model’s ability to classify the three brain subviews, resulting in 2949 test images. As SonoNet lacks dedicated classes for both cervix and transthalamic, we mapped the ”other” class in SonoNet to cervix and transthalamic for the first and second evaluations, respectively. In addition, the performance of the SonoNet models varied across sizes, with SN16, SN32, and SN64 achieving F1 scores of 67.5%, 69.9%, and 67.2%, respectively, where the mid-sized model (SN32) demonstrated the best performance.
@@ 5 Data Availability
This study involved two publicly available datasets for evaluating models on a subset of tasks, as well as private datasets for pretraining and additional evaluations. The public datasets were obtained from PlanesDB (https://zenodo.org/records/3904280) and HC18 (https://zenodo.org/records/1327317). Due to regulations and privacy constraints, the private datasets cannot be released to the public.
@@ 7 Acknowledgements
We express our gratitude to Corniche Hospital in Abu Dhabi for providing prenatal scan data along with fetal heart scans, and to the Department of Health (DOH) Abu Dhabi for their support in approving the study which facilitates access to the anonymous data for internal purposes. We thank Alfred Z. Abuhamad for allowing us to leverage his book for foundation model pretraining. We also thank GE Healthcare for providing data and annotations used for downstream segmentation tasks in the 4-chamber and abdomen views.